# Topologically protected flatness in chiral moiré heterostructures

### v1

@@ Abstract
The observation of delicate correlated phases in twisted heterostructures of graphene and transition metal dichalcogenides suggests an inherent resilience of moiré flat bands against certain types of disorder. We investigate the robustness of moiré flat bands in the chiral limit of the Bistrizer-MacDonald model, applicable to both platforms in certain limits. We show a drastic difference between the protection of the first magic angle and higher magic angles to chiral symmetric disorder such as random higher moiré potential harmonics arising, for instance, from lattice relaxation. We find that the first magic angle is topologically protected by a topological index theorem, similar to the protection of the zeroth Landau level of Dirac fermions, whose flatness withstands any chiral symmetric perturbation such as non-uniform magnetic fields. Focusing on the first magic angle of twisted bilayer graphene, our analysis reveals a hidden non-local constant of motion that permits the decomposition of the non-abelian gauge field induced by inter-layer tunnelings into two decoupled abelian ones, underscoring a topological mechanism for band flatness. Through numerical simulations, we further show the strikingly different robustness of flat bands across protected (first) and unprotected (higher) magic angles in the presence of various types of disorder and identify the scattering processes that are enhanced or suppressed in the chiral limit. Interestingly, we find that the suppression of disorder broadening persists away from the chiral limit and is further accentuated by isolating a single sublattice polarized flat band in energy. Our analysis suggests the Berry curvature hotspot at the top of the K and K’ valence band in the transition metal dichalcogenide monolayers is essential for the stability of its moiré flat bands and their correlated states.
@@ I Introduction
The recent observation of the fractional quantum anomalous Hall effect in twisted monolayers of $\rm MoTe_{2}$ marks the first unequivocal evidence of topologically ordered states in quantum material without external magnetic fields [1, 2, 3, 4]. It paves the way for realizing non-abelian states of matter through superconducting coupling [5, 6, 7, 8, 9, 10, 11, 12, 13, 14], which are essential for the development of topologically protected quantum computing [15]. However, topologically ordered states are typically fragile to perturbations. The original fractional quantum Hall effect in $\rm GaAs$ was only observed for extremely high mobility samples [16]. In comparison, transition metal dichalcogenides (TMDs), such as $\rm MoTe_{2}$, are not necessarily clean monolayers. Their mobility usually lies two orders of magnitude below that of typical samples of $\rm GaAs$ that host fractional quantum Hall states [17, 18, 19, 20, 21]. Additionally, strain, lattice relaxation, and twist angle variations are expected to further disorder the heterostructures built from these monolayers [22, 23, 24, 25, 26]. Nonetheless, fractional Chern insulators [27, 28, 29] were observed in both twisted $\rm MoTe_{2}$ homobilayers and twisted bilayer graphene (TBG) at finite field [30]. This unexpected observation hints at the possibility that these heterostructures carry some intrinsic protection to a disorder that enables the realization of such delicate, strongly correlated phases.
@@
With an external magnetic field, which is typically homogeneous, the additional protection due to the chiral anomaly has limited consequences. Only a slight broadening difference is seen in experiment between the zeroth and higher Landau levels in graphene [39]. In fact, in Hall systems, the protection by a chiral anomaly has little physical importance because the Landau level spread to cyclotron gap ratio can be efficiently reduced by going to larger magnetic fields. Contrast this with moiré materials, where there is no external magnetic field but rather a fictitious gauge field that arises from highly non-uniform interlayer couplings and where the spectral gap around the band of interest cannot be arbitrarily increased using an external experimental knob, and the protection provided by the index theorem can play a crucial role.
@@
In this paper, we explore the protection of flat bands in moiré bilayers endowed from the chiral anomaly in the chiral limit of the Bistrizer-MacDonald model of TBG [40, 41], which has also been shown to be an insightful limit for twisted TMDs upon adding a mass to the Dirac theory of each monolayer [42], and may also apply to other types of heterostructures [43, 44, 45, 46, 47]. This limit permits a formal map between the flat bands at the magic angle and Landau levels under a pseudo, non-abelian gauge field $\vec{\mathcal{A}}$. Naively, we should expect that these states are not protected by the index theorem, which is only applicable when the gauge field is abelian (Sec. II). Remarkably, we discover that for large angles, including the first magic angle, there is a hidden non-local symmetry that allows for the decomposition of the $SU(2)$ gauge field into two decoupled abelian components $U(1)\times U(1)$ (Sec. III). Applied specifically to the chiral limits of TBG and twisted $\rm MoTe_{2}$, this large-angle abelian regime corresponds to twist angles $\theta\gtrsim 0.7^{\circ}$ and $\theta\gtrsim 2.1^{\circ}$, respectively. In the decoupled basis, electrons feel a non-uniform and periodic magnetic field $B^{\prime}$. The matching of the magnetic length introduced by $B^{\prime}$ and the moiré length approximates well the magic condition, as anticipated in Ref. [48] where it was understood as a consequence of the topological Atiyah-Singer index theorem [49, 50]. This leads to striking qualitative differences in the robustness of the different magic angles against chiral preserving disorder that we numerically investigate in Sec. IV by introducing random higher harmonics of the moiré interlayer tunneling potential. We note that our analysis captures the effect of lattice corrugations and strain introduced at the moiré unit cell edges, which is particularly important at small twist angles. We find that the first magic angle is mostly undisturbed by higher harmonics, which makes it possible to consistently define the same magic angle condition throughout the sample despite small deformations at sub-moiré length scales. Importantly, this is not true for higher magic angles, suggesting that it may be impossible to observe higher magic angles in realistic experimental settings, which agrees with asymptotic calculations [51].
@@ II Index theorem
A highly counter-intuitive property of the zero-energy Landau level of Dirac particles in a magnetic field was simultaneously discovered by Aharonov and Casher [31] and Dubrovin and Novikov (who considered the square of the Dirac Hamiltonian) [38]: it remains perfectly flat and extensively degenerate for arbitrary non-uniform magnetic fields provided the total magnetic flux threading the system is a constant. This strikingly contrasts the behavior of all higher Landau levels of the same system and those stemming from quadratic bands, which in similar settings acquire a dispersion proportional to the non-uniformity of the magnetic field.
@@
The topological nature of the AS theorem is clear: it only depends on the integrated magnetic flux but not on the local details and variation of $\mathcal{B}$. The right-hand side of Eq. 2 is nothing but the Chern number of the electromagnetic connection, or gauge field, with which the reader may be more familiar in the context of Berry connection/curvature [52] (to avoid confusion, we stress that these two Chern numbers are a priori unrelated and that the “topology” of the AS theorem is distinct from to band topology). The most direct physical consequence of the AS theorem is that the number of zero modes of $\mathcal{H}_{\rm D}$ is independent of local variations of the gauge potential $\vec{\mathcal{A}}$ and of the metric of the Dirac operator. This is equivalently stated: the density of states at zero energy is topologically protected against all perturbations of the Hamiltonian anti-commuting with the chiral operator. In the case of graphene in a magnetic field, these perturbations physically correspond to strain and lattice corrugations [37, 53].
The AS index also protects the zeroth Landau level of systems with quadratic dispersion provided they can be written down as the square of $\mathcal{H}_{\rm D}$ up to an overall sublattice dependent uniform potential [38, 54]. For particles in a magnetic field, this condition is equivalent to having a $g-$factor exactly equal to 2. In other contexts, it can lead to spin-orbit-like corrections to the quadratic Hamiltonian [42].
@@ III Application to chiral TBG
In this section, we introduce the chiral model for TBG and prove that, for sufficiently large angles, it can be analytically decoupled into two independent Dirac Hamiltonians of the form of Eq. 1 coupled to abelian gauge fields. The flat bands appearing in this abelian regime correspond to the first (experimentally observed) magic angle, whose existence can be understood as a consequence of the AS theorem. We contrast this behavior with higher (theoretically predicted) magic angles, which we show are not topologically protected by the index theorem in Eq. 2 due to their nodal structure. This difference has important consequences regarding the robustness of the flat bands, investigated in Sec. IV.
@@ III.1 Interlayer tunneling as a fictitious gauge field
with $\kappa_{n}$ the $32\pi n/$ 3-rotation of the $\kappa$ corner of the moiré Brillouin zone aligned with the negative $y$-axis, $|\kappa|=2|K|\sin(\theta/2)$, $v$ the Fermi velocity of graphene, and where $\alpha$ measures the strength of $AB$ tunneling $w$ in unit of the typical moiré kinetic energy $|v\kappa|$. The Hamiltonian in the other valley is obtained by time-reversal conjugation.
The suggestive form of Eq. 3 makes it clear that $AB$ interlayer tunnelings in chiral TBG act as a fictitious gauge potential for the two coupled Dirac cones with chiral symmetry $\sigma_{3}$ [55]. Furthermore, the chiral model Eq. 3 is known to host perfectly flat bands at zero energy for $\alpha=0.586,2.221,3.751,\cdots$ [56]. In that context, it is natural to wonder whether these exact flat bands are topologically protected by the index theorem of Eq. 2 or not [57, 48]. At first sight, the answer to this question seems to be negative. Indeed, $\mathcal{A}_{z}$ physically comes from inter-layer tunneling and is entirely carried by off-diagonal layer Pauli matrices. As a result of the trace in Eq. 2, the index corresponding to this gauge field must vanish. In fact, this is a generic feature of SU($N$) gauge fields with $N>1$, whose generators are all traceless.
Figure 1: The three components of the SU(2) fictitious magnetic field strength $\mathcal{B}=B_{1}\mu_{1}+B_{2}\mu_{2}+B_{3}\mu_{3}$ induced by inter-layer tunneling for a small $\alpha=0.2$ (a), the first magic angle $\alpha=0.586$ (b), and the second magic angle $\alpha=3.751$ (c). The different $\alpha$-scaling of the abelian $B_{j=1,2}=\alpha\vec{\nabla}\times\vec{A}^{(j)}$ and non-abelian $B_{3}=2\alpha^{2}\vec{A}^{(1)}\times\vec{A}^{(2)}$ components causes a transition from a weakly non-abelian regime ($\alpha\lesssim 1$) to a deeply non-abelian one ($\alpha\gtrsim 1$). The grey hexagon shows the $\sqrt{3}\times\sqrt{3}$ enlarged moiré unit cell over which the tunneling terms are periodic.
We notice that the fictitious gauge field is almost abelian for small values of $\alpha$ (see Fig. 1). This becomes clear if we split the fictitious field strength $\mathcal{B}=\mathcal{B}_{\rm ab}+\mathcal{B}_{\rm nab}$ into abelian $\mathcal{B}_{\rm ab}$ and non-abelian $\mathcal{B}_{\rm nab}$ components, which respectively read
$\begin{split}\mathcal{B}_{\rm ab}&=B_{1}\mu_{1}+B_{2}\mu_{2},\quad B_{j=1,2}=\alpha\vec{\nabla}\times\vec{A}^{(j)},\\ \mathcal{B}_{\rm nab}&=B_{3}\mu_{3},\quad B_{3}=2\alpha^{2}\vec{A}^{(1)}\times\vec{A}^{(2)}.\end{split}$ (4)
One of the crucial points of this paper is that the abelian and non-abelian parts of $\mathcal{B}$ do not scale with the same power of $\alpha$. As a result, the system is almost abelian for small $\alpha$, i.e. for large twist angles, while it is almost entirely non-abelian for large $\alpha$, i.e. small angles. We illustrate this difference in Fig. 1, where we plot the magnetic strength $B_{1,2,3}$ for a small $\alpha=0.2$, and at the two first magic angles $\alpha=0.586$ and $3.751$.
The weakly non-abelian case, for $\alpha\ll 1$, bears similarity with that of electrons evolving in an underlying non-colinear spin texture producing an effective $SU(2)$ gauge field [58, 59]. In the adiabatic limit where the spin texture slowly varies in space, the system split into two decoupled components, respectively aligned and anti-aligned with the underlying texture, that each experience an abelian gauge potential [60, 61, 62]. Similar adiabatic approximations have been used in different moiré systems to approximate the inter-layer tunneling as an effective $U(1)$ abelian gauge field [63, 64]. For a quantitative description of the physics, the domain of validity of this adiabatic approximation should be identified, and all the correcting terms should be explicitly included.
The near-abelian gauge field observed for small $\alpha$ in Fig. 1 hints at a similar decoupling, thereby separating two distinct regimes in the physics of TBG: an abelian one for $\alpha\lesssim 1$ in which the AS index theorem can topologically protect a flat band, and a non-abelian one for $\alpha\gtrsim 1$ where no topological protection occurs and flat bands are accidental.
@@ III.2 Hidden constant of motion
We now make the distinction between the abelian and non-abelian regime analytically precise by proving that an invertible transformation can decouple the Hamiltonian of chiral TBG (Eq. 3) into two independent abelian $U(1)$ Dirac operators, for which the AS index theorem applies, so long as $\alpha<0.89$.
Our goal is to find the $\alpha$ for which there exists an invertible transformation $U$, making the Dirac operator abelian or diagonal; in other words, we want

### v2

# thanks: These two authors contributed equallythanks: These two authors contributed equally

@@ Abstract
The observation of delicate correlated phases in twisted heterostructures of graphene and transition metal dichalcogenides suggests that moiré flat bands are intrinsically resilient against certain types of disorder. Here, we investigate the robustness of moiré flat bands in the chiral limit of the Bistrizer-MacDonald model – applicable to both platforms in certain limits – and demonstrate drastic differences between the first magic angle and higher magic angles in response to chiral symmetric disorder that arise, for instance, from lattice relaxation. We understand these differences using a hidden constant of motion that permits the decomposition of the non-abelian gauge field induced by interlayer tunnelings into two decoupled abelian ones. At all magic angles, the resulting effective magnetic field splits into an anomalous contribution and a fluctuating part. The anomalous field maps the moiré flat bands onto a zeroth Dirac Landau level, whose flatness withstands any chiral symmetric perturbation such as non-uniform magnetic fields due to a topological index theorem – thereby underscoring a topological mechanism for band flatness. Only the first magic angle can fully harness this topological protection due to its weak fluctuating magnetic field. In higher magic angles, the amplitude of fluctuations largely exceeds the anomalous contribution, which we find results in a physically meaningless chiral operator, and an extremely large sensitivity to microscopic details and an exponential collapse of the single particle gap. Through numerical simulations, we further study various types of disorder and identify the scattering processes that are enhanced or suppressed in the chiral limit. Interestingly, we find that the topological suppression of disorder broadening persists away from the chiral limit and is further accentuated by isolating a single sublattice polarized flat band in energy. Our analysis suggests the Berry curvature hotspot at the top of the $K$ and $K^{\prime}$ valence band in the transition metal dichalcogenide monolayers is essential for the stability of its moiré flat bands and their correlated states.
@@ I Introduction
The recent observation of the fractional quantum anomalous Hall effect in twisted monolayers of $\rm MoTe_{2}$ marks the first unequivocal evidence of topologically ordered states in quantum material without external magnetic fields [1, 2, 3, 4]. It paves the way for realizing non-abelian states of matter [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], which are essential for the development of topologically protected quantum computing [16]. However, topologically ordered states are typically fragile to perturbations. The original fractional quantum Hall effect in $\rm GaAs$ was only observed for extremely high mobility samples [17]. In comparison, transition metal dichalcogenides (TMDs), such as $\rm MoTe_{2}$, are not necessarily clean monolayers. Their mobility usually lies two orders of magnitude below that of typical samples of $\rm GaAs$ that host fractional quantum Hall states [18, 19, 20, 21, 22]. Additionally, strain, lattice relaxation, and twist angle variations are expected to further disorder the heterostructures built from these monolayers [23, 24, 25, 26, 27]. Nonetheless, fractional Chern insulators [28, 29, 30, 31] were observed in both twisted $\rm MoTe_{2}$ homobilayers and twisted bilayer graphene (TBG) at finite field [32]. This unexpected observation hints at the possibility that these heterostructures carry some intrinsic protection to a disorder that enables the realization of such delicate, strongly correlated phases.
@@
With an external magnetic field, which is typically homogeneous, the additional protection due to the chiral anomaly has limited consequences. Only a slight broadening difference is seen in experiment between the zeroth and higher Landau levels in graphene [41]. In fact, in Hall systems, protection by a chiral anomaly has little physical importance because the Landau level spread to cyclotron gap ratio can be efficiently reduced by using larger magnetic fields. In contrast, moiré materials do not rely on an external magnetic field. Instead, they feature a fictitious gauge field arising from non-uniform interlayer couplings. The same couplings are responsible for the spectral gaps around the band of interest, which cannot be arbitrarily tuned using external fields. As a result, the protection offered by the index theorem can play a crucial role.
@@
In this paper, we explore the protection of flat bands in moiré bilayers endowed from the chiral anomaly in the chiral limit of the Bistrizer-MacDonald model of TBG [42, 43], which has also been shown to be an insightful limit for twisted TMDs upon adding a mass to the Dirac theory of each monolayer [44], and may also apply to other types of heterostructures [45, 46, 47, 48, 49, 50, 51]. This limit permits a formal map between the flat bands at the magic angle and Landau levels under a pseudo, non-abelian gauge field $\vec{\mathcal{A}}$. Naively, we should not expect protection by the index theorem, which only applies when the gauge field is abelian (Sec. II). Yet, we discover a hidden non-local symmetry that allows for the decomposition of the $SU(2)$ non-abelian gauge field into two decoupled abelian components $U(1)\times U(1)$ (Sec. III). In the decoupled basis, electrons feel a non-uniform and periodic magnetic field $B^{\prime}$. The matching of the magnetic length introduced by this effective magnetic field and the moiré length provides an analytical criterion for magic angles, as anticipated in Ref. [52] where it was understood as a consequence of the topological Atiyah-Singer index theorem [53, 54]. For all but the largest magic angle however, the fluctuating part of the magnetic field is pathological, and the system cannot fully harness the topological protection offered by the index theorem for reasons we detail in Sec. IV. First, the index theorem only protects the system with respect to chiral symmetric disorder. The large and short-scale oscillations of the effective magnetic field obtained at the higher magic angles corresponds to a quickly oscillating chiral operator, and, in turn, to a very fine-tuned set of disorder realizations keeping the flat bands perfectly flat. In contrast, the chiral operator at the first magic angle is mostly carried on low harmonics of the moiré lattice, with which realistic disorder pattern have high overlap, yielding an effective protection of the band flatness. Second, fluctuations of the effective magnetic field $B^{\prime}$ in higher magic angle are much greater than its average, which is known to yield an exponential squeezing of the spectrum around zero energy [55, 56, 57, 58]. As a result, non-chiral disorder will affect higher magic angles more strongly than the first. Altogether, this leads to striking qualitative differences in the robustness of the different magic angles against chiral preserving disorder that we numerically investigate by introducing random higher harmonics of the moiré interlayer tunneling potential. The first magic angle is mostly undisturbed by higher harmonics, which makes it possible to consistently define the same magic angle condition throughout the sample despite small deformations. This is not true for higher magic angles, suggesting that it is more difficult to observe fragile correlated phases in higher magic angles in realistic experimental settings. In Sec. V, we consider the case of twisted TMDs where the anomaly applies thanks to the Berry curvature hotspots at the top of the $K$ and $K^{\prime}$ valence band.
@@ II Index theorem
A highly counter-intuitive property of the zero-energy Landau level of Dirac particles in a magnetic field was simultaneously discovered by Aharonov and Casher [33] and Dubrovin and Novikov (who considered the square of the Dirac Hamiltonian) [40]: it remains perfectly flat and extensively degenerate for arbitrary non-uniform magnetic fields provided the total magnetic flux threading the system is a constant. This behavior strikingly contrasts with that of all higher Landau levels of the same system and those stemming from quadratic bands, which in similar settings acquire a dispersion proportional to the magnetic field’s non-uniformity.
@@
The topological nature of the AS theorem is clear: it only depends on the integrated magnetic flux but not on the local details and variation of $\mathcal{B}$. The right-hand side of Eq. 2 is nothing but the Chern number of the electromagnetic connection, or gauge field, with which the reader may be more familiar in the context of Berry connection/curvature [59] (to avoid confusion, we stress that these two Chern numbers are a priori unrelated and that the “topology” of the AS theorem is distinct from to band topology). The most direct physical consequence of the AS theorem is that the number of zero modes of $\mathcal{H}_{\rm D}$ is independent of local variations of the gauge potential $\vec{\mathcal{A}}$ and of the metric of the Dirac operator. Put differently, the density of states at zero energy is topologically protected against all perturbations of the Hamiltonian anticommuting with the chiral operator. In the case of graphene in a magnetic field, these perturbations physically correspond to strain and lattice corrugations [39, 60, 61].
The AS index also protects the zeroth Landau level of systems with quadratic dispersion provided they can be written down as the square of $\mathcal{H}_{\rm D}$ up to an overall sublattice dependent uniform potential [40, 62]. This condition is equivalent to having a $g$-factor precisely equal to 2. In some contexts, it can lead to spin-orbit-like corrections to the quadratic Hamiltonian [44].
While the index theorem enforces the existence of an extensive number of zero modes, there is a cost to pay for this flatness when fluctuations of the magnetic field are large compared to its average: the spectral gap separating the zero modes from the rest of the spectrum become exponentially small in the amplitude of magnetic field fluctuations. This exponential squeezing of the full spectrum as a function of the maximum amplitude of the effective gauge field $\mathcal{A}$ was proven for the specific case of the chiral Hamiltonian of TBG in Ref. [55], and follows from very generic considerations on Landau operators [56, 57, 58]. Intuitively, this scaling of the gap comes from the fact that the existence of zero modes requires a discrete spectrum of the Hamiltonian around which the resolvent is large [63], allowing to form eigenvectors with energy $\delta E$ away from the zero modes for perturbations of the Hamiltonian that are exponentially small in $\delta E$.
In realistic settings, disorder does not solely occur in the chiral channel. It nevertheless carries a chiral component, against which the system is protected, and an non-chiral one. When the chiral symmetry varies slowly on the moiré scale, the harmless chiral part will be large as the most important contribution of the disorder carried by the low moiré harmonics strongly overlap with the chiral protected channels. This is what happens at the first magic angle. In contrast, when the chiral operator wildly fluctuates within the moiré unit cell, the overlap of any realistic disorder realization will have no overlap with it and be almost entirely unprotected. This is what we observe at the higher magic angle, which we discuss in detail in Sec. III and IV.
@@ III Application to chiral TBG
In this section, we introduce the chiral model for TBG and prove that it decouples into two independent Dirac Hamiltonians akin to Eq. 1 coupled to abelian gauge fields. In this abelian basis, flat bands appear when the total magnetic field flux per moiré unit cell equals $2\pi$ due to the AS theorem, which provides an intuitive picture for the emergence of the magic angles. The magnetic field distribution in the unit cell at the different magic angles will be crucial to understanding the robustness of the flat bands against disorder, investigated in Sec. IV.
@@ III.1 Interlayer tunneling as a fictitious gauge field
with $\kappa_{n}$ the $32\pi n/$ 3-rotation of the $\kappa$ corner of the moiré Brillouin zone aligned with the negative $y$-axis, $|\kappa|=2|K|\sin(\theta/2)$, $v$ the Fermi velocity of graphene, and where $\alpha$ measures the strength of $AB$ tunneling $w$ in unit of the typical moiré kinetic energy $|v\kappa|$. The Hamiltonian in the other valley is obtained by time-reversal conjugation. The system has $C_{3z}$, $C_{2x}$, $C_{2y}T$ and $C_{2z}T$ symmetries, with $T$ the time-reversal operator [64].
The suggestive form of Eq. 3 makes it clear that $AB$ interlayer tunnelings in chiral TBG act as a fictitious gauge potential for the two coupled Dirac cones with a chiral anticommuting symmetry $\sigma_{3}$ [65]. Furthermore, the chiral model Eq. 3 is known to host perfectly flat bands at zero energy for $\alpha=0.586,2.221,3.751,\cdots$ [55, 66]. In that context, it is natural to wonder whether these exact flat bands are topologically protected by the index theorem of Eq. 2 or not [67, 52]. At first sight, the answer to this question seems to be negative. Indeed, $\mathcal{A}_{z}$ physically comes from inter-layer tunneling and is entirely carried by off-diagonal layer Pauli matrices. As a result of the trace in Eq. 2, the index corresponding to this gauge field must vanish. In fact, this is a generic feature of SU($N$) gauge fields with $N>1$, whose generators are all traceless.
@@ III.2 Hidden constant of motion
We overcome this problem by proving that an invertible transformation can decouple the Hamiltonian of chiral TBG (Eq. 3) into two independent abelian $U(1)$ Dirac operators, for which the AS index theorem may apply. Our goal is now to find an invertible transformation $U$ making the Dirac operator abelian or diagonal; in other words, we want